Question 3.

A hypothetical economy produces only four types of goods domestically; food, clothing, housing, machines. The quantities produced and their respective prices are shown below for two different years, namely 2000 (base year) and 2002 (current year)

Q(2000) P(2000) Q(2002) P(2002)
Food 12 10 15 11
Clothing 15 15 18 18
Housing 14 20 20 28
Machines 20 25 25 28

A. Calculate the inflation rate over the period 2000-2002 based on the CPI measures of the price level.

B. Explain the why CPI measure of inflation overestimate the actual growth in consumer prices by about one percent.
Answer #1

a) For finding the CPI and inflation rate, find the cost of basket of goods using base year prices and current year prices. This can be done by finding the products of price and quantity for all goods and adding them up.

Q(2000) P(2000) Q(2002) P(2002) Cost of basket 2000 Cost of basket 2002 at base year level
Food 12 10 15 11 120 132
Clothing 15 15 18 18 225 270
Housing 14 20 20 28 280 392
Machines 20 25 25 28 500 560
Total cost 1125 1354

The last two columns contain information regarding CPI. CPI index for base year is 100 because we have

CPI = Cost of basket in current year (2000)*100/Cost of basket in current year (2000) = 100

Now CPI in current year 2002 = Cost of basket in current year (2002)*100/Cost of basket in current year (2000) = 1354*100/1125 = 120.36

Find the inflation rate using CPI, inflation rate = (CPI 2002 - CPI 2000)*100/CPI 2000 = (120.36 - 100)*100/100 = 20.36% This is the rate of inflation

b) CPI usually overstates the inflation rate since it suffers from many biases like the substitution bias. It measures the changes in prices only and not in the level of production so actual inflation is generally less.